### v1

# Dynamical Mean-Field Theory of Complex Systems on Sparse Directed Networks

Abstract
Although real-world complex systems typically interact through sparse and heterogeneous networks, analytic solutions of their dynamics are limited to models with all-to-all interactions. Here, we solve the dynamics of a broad range of nonlinear models of complex systems on sparse directed networks with a random structure. By generalizing dynamical mean-field theory to sparse systems, we derive an exact equation for the path-probability describing the effective dynamics of a single degree of freedom. Our general solution applies to key models in the study of neural networks, ecosystems, epidemic spreading, and synchronization. Using the population dynamics algorithm, we solve the path-probability equation to determine the phase diagram of a seminal neural network model in the sparse regime, showing that this model undergoes a transition from a fixed-point phase to chaos as a function of the network topology.
Introduction.
Complex dynamical systems are modeled by $N$ degrees of freedom $x_{i}(t)$ ($i=1,\dots,N$) that evolve in time according to the coupled differential equations
$\dot{x}_{i}(t)=-f(x_{i})+\sum_{j=1}^{N}A_{ij}g(x_{i},x_{j})+\xi_{i}(t),$ (1)
where $\{A_{ij}\}_{i,j=1,\dots,N}$ defines the interaction network and $\xi_{i}(t)$ is an uncorrelated Gaussian noise, with mean zero and variance $\sigma^{2}$. The variable $\xi_{i}(t)$ models external random perturbations, the function $f(x)$ governs the dynamics in the absence of interactions, and the kernel $g(x,x^{\prime})$ shapes the pairwise couplings. Equation (1) models the nonequilibrium dynamics of neural networks Sompolinsky et al. (1988); Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Martorell et al. (2024) and ecosystems Opper and Diederich (1992); Bunin (2017); Altieri et al. (2021), epidemic spreading Mata and Ferreira (2013); Mieghem (2012); Goltsev et al. (2012); Pastor-Satorras et al. (2015), synchronization phenomena Stiller and Radons (1998); Rodrigues et al. (2016); Prüser et al. (2024), opinion dynamics Baumann et al. (2020, 2021), and multivariate Ornstein-Uhlenbeck processes Godrèche and Luck (2018); Crisanti and Paoluzzi (2023); Ferraro et al. (2024). Table 1 specifies $f(x)$ and $g(x,x^{\prime})$ for paradigmatic models of complex behavior.
Model $f(x)$ $g(x,x^{\prime})$
Ornstein-Uhlenbeck (OU) process Godrèche and Luck (2018) $x$ $x^{\prime}$
SIS model of epidemic spreading Mata and Ferreira (2013) $x$ $(1-x)x^{\prime}$
Lotka-Volterra (LV) model Bunin (2017); Altieri et al. (2021) $\;\,x(x-1)$ $xx^{\prime}$
Neural network (NN) model Sompolinsky et al. (1988); Crisanti and Sompolinsky (2018) $x$ $\tanh(x^{\prime})$
Kuramoto model Rodrigues et al. (2016) 0 $\sin(x^{\prime}-x)$
Table 1: Explicit form of $f(x)$ and $g(x,x^{\prime})$ in complex systems modeled by Eq. (1).
The foremost problem in the study of complex systems is how to reduce the dynamics of many interacting elements to the dynamics of a few variables Thibeault et al. (2024). Dynamical mean-field theory (DMFT) Martin et al. (1973); Coolen (2001); Hertz et al. (2016) is a powerful method to tackle this problem in the limit $N\rightarrow\infty$, yielding a solution in terms of the path-probability for the effective dynamics of a single degree of freedom. The application of DMFT to models described by Eq. (1) has been attracting an enormous interest Kadmon and Sompolinsky (2015); Martí et al. (2018); Crisanti and Sompolinsky (2018); Roy et al. (2019); Martorell et al. (2024); Bunin (2017); Galla (2018); Schuecker et al. (2018); Altieri et al. (2021); Prüser et al. (2024); Crisanti and Paoluzzi (2023); Poley et al. (2023); Ferraro et al. (2024); Aguirre-López (2024); Park et al. (2024); Pham and Kaneko (2024), especially in the context of neural networks and ecosystems. Phase diagrams of these models reveal a rich phenomenology, including different types of phase transitions Bunin (2017); Martorell et al. (2024), chaotic behavior Sompolinsky et al. (1988); Martí et al. (2018), and coexistence of multiple attractors Bunin (2017); Altieri et al. (2021); Ros et al. (2023). Despite this substantial theoretical progress, DMFT is currently limited to fully-connected networks, where each dynamical variable is coupled to all others by means of Gaussian interaction strengths. Nevertheless, the interactions in real-world complex systems are known to be sparse and heterogeneous Newman (2010); Guimarães (2020). Sparseness indicates that each element of the system interacts on average with a finite number of others, while heterogeneity refers to fluctuations in the local topology of the interaction network. How to integrate these features in the formalism of DMFT for systems modeled by Eq. (1) remains an unresolved challenge, and even basic questions, such as characterizing the phase diagram of sparse complex systems, are still out of reach.
On the other hand, the dynamics of complex systems on sparse networks has been extensively studied in the case of Ising spins Hatchett et al. (2004); Neri and Bollé (2009); Mimura and Coolen (2009); Roudi and Hertz (2011); Aurell and Mahmoudi (2012); Del Ferraro and Aurell (2015); Aurell et al. (2017); Vázquez et al. (2017); Behrens et al. (2023, 2024). In this context, both the cavity method Neri and Bollé (2009) and DMFT Hatchett et al. (2004); Mimura and Coolen (2009) provide an analytic solution in terms of the path-probability for the effective dynamics of a single unit. However, the presence of symmetric interactions induce a temporal feedback, which, in the particular case of sparse systems, leads to an exponential growth of the path-probability dimension Hatchett et al. (2004); Mimura and Coolen (2009); Behrens et al. (2023), rendering numerical computations unfeasible. When the interactions are directed or unidirectional, this issue does not arise, allowing the path-probability equation to be efficiently solved Derrida et al. (1987); Hatchett et al. (2004); Neri and Bollé (2009).
Inspired by these results for Ising spins, in this letter we solve the dynamics of models governed by Eq. (1) on sparse directed networks with an heterogeneous topology. The solution represents a foundational step in the study of sparse complex systems, as it ultimately incorporates the role of the network structure in the formalism of DMFT. Besides that, directed networks are interesting from a practical viewpoint, because they model the non-reciprocal interactions in real-world complex systems Asllani et al. (2018), including the synaptic connections in the human cortex Peng et al. (2024) and the trophic interactions in food-webs Dunne et al. (2002); Bascompte (2009); Guimarães (2020).
By generalizing DMFT to sparse systems, we obtain an exact equation for the path-probability describing the effective dynamics of a single variable for $N\rightarrow\infty$. We solve this equation for different models in table 1 by using the population dynamics algorithm Mézard and Parisi (2001, 2003), which is a fundamental tool in the study of sparse disordered systems Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). The excellent agreement between our theoretical results and numerical simulations of finite systems confirms that our solution is valid for several models.
As an application, we determine the phase diagram of the neural network model by Sompolinsky et al Sompolinsky et al. (1988) in the sparse regime. We show that the phase diagram displays a trivial fixed-point phase, a nontrivial fixed-point phase, and a chaotic region. Our results for the transition lines, based on the dynamics of macroscopic observables, are consistent with the universal critical lines derived from random matrix theory Neri and Metz (2016, 2020); Metz and Neri (2021). Remarkably, we provide evidence that the transition from the nontrivial fixed-point phase to the chaotic regime coincides with the vanishing of the gap between the leading eigenvalue and the bulk of the network spectrum.
Sparse directed networks.
Let $A_{ij}=C_{ij}J_{ij}$ be the elements of the $N\times N$ interaction matrix $\boldsymbol{A}$. The binary variables $C_{ij}\in\{0,1\}$ determine the network topology, while $J_{ij}\in\mathbb{R}$ controls the pairwise interaction strengths. If $C_{ij}=1$, there is a directed edge $j\rightarrow i$ pointing from node $j$ to $i$, while $C_{ij}=0$ otherwise. We also set $C_{ii}=0$. The indegree $K_{i}=\sum_{j=1}^{N}C_{ij}$ and the outdegree $L_{i}=\sum_{j=1}^{N}C_{ji}$ count the number of links entering and leaving node $i$ Newman (2010), respectively. The random variables $\{C_{ij}\}_{i\neq j}$ follow the distribution
($\mathbb{P}(\{C_{ij}\})$ $=\frac{1}{\mathcal{N}}\prod_{i\neq j=1}^{N}\left[\frac{c}{N}\delta_{C_{ij},1}+\left(1-\frac{c}{N}\right)\delta_{C_{ij},0}\right]$
×$\times\prod_{i=1}^{N}\delta_{K_{i},\sum_{j=1}^{N}C_{ij}}\delta_{L_{i},\sum_{j=1}^{N}C_{ji}},$ (2)
where $\mathcal{N}$ is the normalization constant. The degrees $\{K_{i},L_{i}\}_{i=1,\dots,N}$ are independent and identically distributed random variables drawn from $p_{k,\ell}=p_{{\rm in},k}\,p_{{\rm out},\ell}$, where $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ are the indegree and the outdegree distribution Newman (2010); Metz et al. (2019), respectively. The network parameter $c$ is the average degree
$c=\sum_{k=0}^{\infty}k\,p_{{\rm in},k}=\sum_{\ell=0}^{\infty}\ell\,p_{{\rm out},\ell}.$ (3)
Equation (2) defines a network ensemble where directed links are randomly placed between pairs of nodes with probability $c/N$, subject to the prescribed degree sequences generated from $p_{k,\ell}$. In the limit $N\rightarrow\infty$, the network samples generated from Eq. (2) are equivalent to those produced by the configuration model Newman et al. (2001); Fosdick et al. (2018). The coupling strengths $\{J_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables drawn from a distribution $p_{J}$ with mean $\mu_{J}$ and variance $\sigma_{J}^{2}$. The distributions $p_{k,\ell}$ and $p_{J}$ fully specify the network ensemble, allowing for a systematic investigation of how network heterogeneities impact the dynamics of complex systems.
Solution through DMFT.
We solve the coupled dynamics of Eq. (1) on directed networks by using dynamical mean-field theory (DMFT) Martin et al. (1973); Coolen (2001); Hertz et al. (2016). We consider the sparse regime, where the mean degree $c$ is finite, independent of $N$. DMFT is based on the generating functional
[…]
of the probability density $\mathcal{P}[\boldsymbol{x}]$ of observing a dynamical path of states $\boldsymbol{x}(t)=(x_{1}(t),\dots,x_{N}(t))$ in a fixed time interval. The correlation functions of $\{x_{i}(t)\}_{i=1,\dots,N}$ follow from the derivatives of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ with respect to the external sources $\boldsymbol{\psi}(t)=(\psi_{1}(t),\dots,\psi_{N}(t))$. In particular, the $n$th-moment of the local variable $x_{i}(t)$,
$\langle x_{i}^{n}(t)\rangle=\int\left(\prod\limits_{i=1}^{N}Dx_{i}\right)x_{i}^{n}(t)\mathcal{P}[\boldsymbol{x}]=(-i)^{n}\frac{\delta^{n}\mathcal{Z}\left[\boldsymbol{\psi}\right]}{\delta\psi_{i}^{n}(t)}\Bigg{|}_{\boldsymbol{\psi}=0},$ (5)
yields the time-evolution of the macroscopic quantities
$m(t)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\langle x_{i}(t)\rangle,\quad q(t)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\langle x_{i}^{2}(t)\rangle.$ (6)
Clearly, the generating functional fulfills $\mathcal{Z}\left[0\right]=1$.
In the supplemental material Metz , we calculate the average of $\mathcal{Z}\left[\boldsymbol{\psi}\right]$ over the network ensemble defined by Eq. (2) for finite $c$, recasting the problem in terms of the solution of a saddle-point integral. More importantly, we give a clear physical interpretation of the order-parameters, simplifying the saddle-point equations and obtaining a feasible solution for the dynamics in the sparse regime. In the limit $N\rightarrow\infty$, the microscopic dynamical variables decouple, and the path-probability $\mathcal{P}[x]$ for the effective dynamics of a single variable $x(t)$ is determined from
$\mathcal{P}[x]=\sum\limits_{k=0}^{\infty}p_{{\rm in},k}\int\left(\prod\limits_{j=1}^{k}Dx_{j}\mathcal{P}[x_{j}]\right)\int\left(\prod\limits_{j=1}^{k}dJ_{j}p_{J}(J_{j})\right)$
×$\times\Bigg{\langle}\delta_{F}\Bigg{[}\dot{x}(t)+f(x(t))-\sum\limits_{j=1}^{k}J_{j}\,g(x(t),x_{j}(t))-\xi(t)\Bigg{]}\Bigg{\rangle}_{\xi},$ (7)
where $\delta_{F}$ is the functional Dirac-$\delta$, and $\langle\dots\rangle_{\xi}$ represents the average over the uncorrelated Gaussian noise $\xi(t)$. The macroscopic observables are given by $m(t)=\langle x(t)\rangle_{*}$ and $q(t)=\langle x^{2}(t)\rangle_{*}$, where
$\langle x^{n}(t)\rangle_{*}=\int Dx\,x^{n}(t)\mathcal{P}[x]\quad(n=1,2)$ (8)
is the average over the effective dynamics governed by $\mathcal{P}[x]$. The self-consistent Eq. (7) is the exact solution of a broad class of models (see table 1) on sparse directed networks with arbitrary distributions $p_{J}$ and $p_{k,\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$. We solve the more general case of networks with correlated indegrees and outdegrees in Metz . The solution of Eq. (7) determines the time-evolution of the full probability distribution of the microscopic variables in the limit $N\rightarrow\infty$.
Figure 1: Dynamics of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ for the SIS model, LV model, and NN model (see table 1) without external noise ($\sigma=0$). Results are shown for directed random networks with average degree $c=5$ and two indegree distributions (Eq. (9)): Poisson (left column) and geometric (right column). The distribution $p_{J}$ of the coupling strengths has mean $\mu_{J}$ and standard deviation $\sigma_{J}=0.1$. For the LV and NN models, $p_{J}$ is Gaussian; for the SIS model, $p_{J}$ is uniform. Solid lines are solutions of Eq. (7) using the population dynamics algorithm with $5\times 10^{4}$ paths. Symbols denote numerical simulations of Eq. (1) for an ensemble of $10$ independent random networks generated from the configuration model with $N=4000$ nodes. Vertical bars are the standard deviation of the macroscopic observables.
From now on, we set $\sigma=0$ in Eq. (7) and study the dynamics in the absence of external noise. Below we present results for Poisson and geometric indegrees, where the indegree distribution $p_{{\rm in},k}$ is given by
$p_{{\rm in},k}=\frac{c^{k}e^{-c}}{k!}\quad{\rm and}\quad p_{{\rm in},k}=\frac{c^{k}}{(c+1)^{k+1}},$ (9)
respectively.
Equation (7) is formally similar to other distributional equations appearing in the study of sparse disordered systems Mézard and Parisi (2001, 2003); Kühn (2008); Metz et al. (2019); Abou-Chacra et al. (1973); Mézard and Montanari (2009). Therefore, we can numerically solve this equation using the population dynamics algorithm Mézard and Parisi (2001); Kühn (2008); Metz et al. (2019). In the standard version of this algorithm Mézard and Parisi (2001); Metz et al. (2019), a probability density is parametrized by a population of stochastic variables. Here, we generalize the algorithm to calculate $\mathcal{P}[x]$ by introducing a population of dynamical trajectories. At each iteration step, a single path is chosen randomly from the population and updated according to the differential equation imposed by the Dirac-$\delta_{F}$ in Eq. (7). After sufficient iterations, the population of paths converges to a stationary distribution, providing a numerical solution for $\mathcal{P}[x]$. A detailed account of the algorithm is in Metz .
In Fig. 1, we compare the solutions of Eq. (7) with numerical simulations of the original coupled dynamics, Eq. (1), for finite $N$. The panels showcase the time-evolution of the mean $m(t)$ and the standard deviation $\sqrt{q(t)-m^{2}(t)}$ of the microscopic variables for different network topologies and three important models of complex systems (see table 1): the NN model of Sompolinsky et al. (1988), the LV model, and the SIS model of epidemic spreading. In all cases, the agreement between our theoretical results for $N\rightarrow\infty$ and numerical simulations for large $N$ is excellent, confirming the exactness of Eq. (7).
An important question is whether Eq. (7) recovers the analytic results of fully-connected models as $c\rightarrow\infty$ Sompolinsky et al. (1988); Crisanti and Sompolinsky (2018); Galla (2018). We show in Metz that this is generally not the case. In the high-connectivity limit $c\rightarrow\infty$, degree fluctuations remain significant, and $\mathcal{P}[x]$ depends on the full distribution $\nu_{\rm in}(\kappa)$ of rescaled indegrees $\kappa_{i}=K_{i}/c$ ($i=1,\dots,N$). The well-known effective dynamics on fully-connected networks is only recovered when $\nu_{\rm in}(\kappa)=\delta(\kappa-1)$. This breakdown in the universality of fully-connected models due to degree fluctuations was anticipated in Metz and Silva (2020); Metz and Peron (2022); Silva and Metz (2022); Ferreira and Metz (2023).
Phase diagram of neural networks.
Now we derive the phase diagram of the NN model on sparse directed networks for $\sigma=0$. In this context, $x_{i}(t)\in\mathbb{R}$ represents the synaptic current at neuron $i$, $f(x)=x$, and $g(x,x^{\prime})=\tanh{\left(x^{\prime}\right)}$. Let us order the complex eigenvalues $\{\lambda_{\alpha}\}_{\alpha=1,\dots,N}$ of $\boldsymbol{A}$ according to their real parts as ${\rm Re}\lambda_{1}>{\rm Re}\lambda_{2}>\dots>{\rm Re}\lambda_{N}$. A linear stability analysis of Eq. (1) shows that $\boldsymbol{x}=0$ is stable if ${\rm Re}\lambda_{1}<1$. Based on previous analytic results for the spectra of directed networks for $N\rightarrow\infty$ Neri and Metz (2016, 2020); Metz and Neri (2021), we find that the trivial fixed-point is stable provided $c<c_{\rm stab}$, where
$c_{\rm stab}=\begin{cases}1/\mu_{J}&\,\,\text{if $c>1+\sigma_{J}^{2}/\mu_{J}^{2}$},\\ 1/(\sigma_{J}^{2}+\mu_{J}^{2})&\text{ if $c\leq 1+\sigma_{J}^{2}/\mu_{J}^{2}$ }.\end{cases}$ (10)
The two distinct regimes in Eq. (10) result from the gap-gapless transition in the spectrum of $\boldsymbol{A}$ Neri and Metz (2016). For $c>1+\sigma_{J}^{2}/\mu_{J}^{2}$, the spectral gap $|\lambda_{1}-\lambda_{2}|$ remains finite as $N\rightarrow\infty$, while it vanishes for $c\leq 1+\sigma_{J}^{2}/\mu_{J}^{2}$. Equation (10) holds for $\mu_{J}>0$ and $c>1$. The later condition ensures that directed random networks contain a giant strongly connected component Timár et al. (2017).
Figure 2: Phase diagram $(c,\sigma_{J})$ of the NN model for fixed $10<\mu_{J}<$ 1. The macroscopic variables $m(t)$ and $q(t)$ relax to trivial and nontrivial fixed-points in phases I and II, respectively. In the chaotic phase III, $m(t)$ shows aperiodic oscillations around zero. The dotted line at $c=1$ marks the percolation threshold, while the dashed curve identifies the gap-gapless transition (see the main text). The inset illustrates the sensitivity of $m(t)$ to small perturbations in the initial conditions. Each colored curve represents the numerical solution of Eq. (7) using population dynamics with $5\times 10^{5}$ paths, with initial states sampled from an uniform distribution with unit mean and standard deviation of $\mathcal{O}(10^{-5})$. The red dot is located at $(c^{*},\sigma^{*}_{J})=(\mu_{J}^{-1},\sqrt{\mu_{J}(1-\mu_{J})})$.
We emphasize that the above stability analysis does not give insights into the model’s dynamical behavior. Therefore, we study the dynamics of $m(t)$ and $q(t)$ by solving Eq. (7). Figure 2 presents the resulting phase diagram. In phase I, $m(t)$ and $q(t)$ relax exponentially to the trivial solution $m=q=0$, while in phase II, $m(t)$ and $q(t)$ evolve to nonzero fixed-points. In phase III, the neural network exhibits chaotic behavior, characterized by slow and aperiodic oscillations of $m(t)$ around zero. The inset in Fig. 2 illustrates the sensitivity of $m(t)$ to small perturbations in the initial conditions, a hallmark of deterministic chaos Baker and Gollub (1996). In phase III, the distribution of neural firing rates $\tanh(x_{i}(t))$ does not evolve to a stationary profile. The phase diagram in Fig. 2 is universal, as the transition lines depend only on the first and second moments of $p_{J}$ and $p_{{\rm in},k}$.
Figure 3: Continuous transition between the fixed-point phases for the NN model on directed random networks. The coupling strengths follow a Gaussian distribution with mean $\mu_{J}=1/3$ and standard deviation $\sigma_{J}=0.1$. The results are derived from Eq. (7) using population dynamics with $5\times 10^{5}$ paths. (a) Fixed-point solution $m(t)=m$ as a function of $c$ for Poisson and geometric indegrees (Eq. (9)). (b) Dynamics of $m(t)$ across the transition for Poisson indegrees.
Figure 3 characterizes the transition between phases I and II. As $c$ approaches the critical mean degree $c^{*}=\mu_{J}^{-1}$ from above, the nontrivial fixed-point $m=m(t)$ vanishes continuously. The critical point $c^{*}$ is independent of $p_{{\rm in},k}$, as predicted by Eq. (10), and $m(t)$ relaxes exponentially fast inside phases I and II. Due to critical slowing down, the numerical solution of Eq. (7) becomes computationally more demanding near $c^{*}$. In phase II, the neuronal firing rates evolve to a stationary distribution with a finite variance.
Figure 4: Dynamics of $m(t)$ and $q(t)$ for the NN model on directed random networks with Poisson indegrees. Random coupling strengths follow a Gaussian distribution with mean $\mu_{J}=1/3$ and standard deviation $\sigma_{J}$. The results are derived from the solutions of Eq. (7) using population dynamics with $5\times 10^{5}$ paths. (a) Transition between phases II and III for $c=5$. The critical value of $\sigma_{J}$ is consistent with the dashed line in Fig. 2. (b) Transition between phases I and III for $\sigma_{J}=0.6$. The critical value of $c$ is consistent with Eq. (10).
Figure 4 shows the dynamics of $m(t)$ and $q(t)$ across the transition between the fixed-point phases and the chaotic regime. Since $m(t)$ remains close to zero in both phases I and III, this is not an effective parameter to probe the transition between them. The second moment $q(t)$, instead, clearly identifies the transition between phases I and III, as seen in Fig. 4(b). In this respect, this is analogous to the transition between the paramagnetic and the spin-glass phase Nishimori (2001); Martorell et al. (2024). The critical line between phases II and III is not determined by the above stability analysis. Figure 4(a) illustrates the dynamics of $m(t)$ as the variance $\sigma_{J}^{2}$ of the couplings increases. Notably, the value of $\sigma_{J}$ marking the sudden drop of $m(t)$ and the onset of chaotic oscillations is consistent with the dashed line in Fig. 2 (see Metz for additional results). This strongly suggests that the transition between phases II and III is governed by the gap $|\lambda_{1}-\lambda_{2}|$ in the spectrum of directed random networks.
Conclusions.
We have developed a dynamical mean-field theory of complex systems on sparse directed networks, deriving an exact path-probability equation for the effective dynamics in the limit $N\rightarrow\infty$. Unlike sparse models with symmetric interactions Hatchett et al. (2004); Neri and Bollé (2009), our path-probability equation can be numerically solved using population dynamics Metz et al. (2019). We confirmed the exactness of our general solution by comparing it with numerical simulations of fundamental models in the study of epidemic spreading, neural networks, and ecosystems. Finally, we applied our solution to determine the complete phase diagram of the sparse and directed version of the canonical neural network model of Sompolinsky et al. (1988).
Our work paves the way for exploring the role of sparse heterogeneous networks on the dynamics of ecosystems, coupled oscillators, epidemic spreading, and beyond. Future works include determining the phase diagram of the sparse Lotka-Volterra model Bunin (2017), the influence of the external noise on phase diagrams Schuecker et al. (2018), and the role of network heterogeneities on the critical exponents of complex systems Ferreira and Metz (2023); de Pirey and Bunin (2024). Equation (7) also forms the basis for calculating correlation functions Martí et al. (2018) and deriving approximate dynamical equations for macroscopic order-parameters.
Acknowledgements.
F. L. M. thanks Tuan Minh Pham for many stimulating discussions. The author also thanks Thomas Peron and Tuan Minh Pham for their valuable comments on the manuscript.